OpenStudy (anonymous):

Solve x2 + 10x + 24 = 0 by completing the square. Show your work for full credit.

OpenStudy (lgbasallote):

would you settle for half credit? lol

OpenStudy (anonymous):

1) subtract 24 from both sides 2) divides 10 by two and then square it to get 25 3) then your equation would be x2+10x+25=24+25 4) your answer after that would be x2+10x+25=49 5) take the square root of the whole eqaution 6) then your answer would be x+5= +- 7 7) x would equal to 2 or -12 Hope that helped you!!!!

OpenStudy (anonymous):

thank you so much.. it did

OpenStudy (anonymous):

your welcome